A grating has 10,000 lines per cm. Calculate the maximum number of principal maxima that can be formed for light of wavelength 487 nm.
1
Expert's answer
2020-02-10T09:44:22-0500

The distance between slits is "\dfrac{1}{10000}=10^{-4}cm=10^{-6}cm"

Once a value for the diffraction grating’s slit spacing d has been determined, the angles for the sharp lines can be found using the equation

"dsin\theta=m\lambda"

where "\theta" is angle between normal and maximum direction; m is number of principal maxima;

λ=487nm is wavelength. Therefore

"m=dsin\theta/\lambda"


Using the properties of sine function

"sin\u03b8\u2a951"


Hence

"m\u2a95 \n\n\u200b \n\\dfrac{d }{\\lambda} =\\dfrac{10^{-6}}{487\\times 10^{-9}}=2.053 \n\n\u200b"


As result get the maximum number of principal maxima is equal to 2.